OpenStudy (anonymous):

I'm doing "Polynomial and Rational Functions" in my College Algebra Course. I need help doing this problem: 76.) f(x)=-x^4 + 6x^3 - 9x^2. The instructions say "Sketch the graph of each function." Anyone willing to solve the problem and explain to me how they did it? Thank you! :)

OpenStudy (jdoe0001):

have you done quadratic factoring yet?

OpenStudy (anonymous):

Yes I have

OpenStudy (jdoe0001):

\(\large { f(x)=-x^4 + 6x^3 - 9x^2\implies \begin{array}{cccllll} -x^2(x^2&-6x&+9)\\ &\uparrow &\uparrow \\ &-3-3&-3\cdot -3 \end{array} }\) any ideas on the zeros or "solutions"?

OpenStudy (anonymous):

Wouldn't you factor it as \[x ^{2}(x ^{2}-4x+3)\] ?

OpenStudy (jdoe0001):

hmmm \((x^2-6x+9)?\) ....

OpenStudy (jdoe0001):

....not sure where the -4x fits in or +3 though

OpenStudy (anonymous):

Wait, never mind. I was looking at the wrong problem in my book :D

OpenStudy (anonymous):

you're correct

OpenStudy (jdoe0001):

so.... what.... do we end up with anyway?

OpenStudy (anonymous):

After factoring, it should be x=0, x=3, and x=3. So, wouldn't the zero multiplicity be 2 and 1?

OpenStudy (anonymous):

No, I think the zero multiplicity is actually 2 because after factoring, it would look like this: \[x ^{2}(x-3)^{2}\]

OpenStudy (jdoe0001):

2 , you're correct

OpenStudy (jdoe0001):

hmmm so... we know that \(\large { f(x)=-x^4 + 6x^3 - 9x^2\implies \begin{array}{cccllll} -x^2(x^2&-6x&+9)\\ &\uparrow &\uparrow \\ &-3-3&-3\cdot -3 \end{array} \\ \quad \\ 0=-x^2(x-3)^2\implies \begin{cases} x=0&2\ multiplicity\\ x=3&2\ multiplicity \end{cases} }\)

OpenStudy (jdoe0001):

when the multiplicity is an EVEN number recall that the graph only makes a U-Turn at the point, it doesn't cross the x-axis so... to know where is UP or DOWN..... pick one value for "x" before and after those u-turn points and graph it away :)

OpenStudy (anonymous):

wouldn't the graph be upside down because the leading coefficient is a negative?

OpenStudy (jdoe0001):

I'd think so, yes

OpenStudy (anonymous):

So what would the graph look like?